OpenStudy (anonymous):

Calculus Area help? use the properties of summation and Theorem 4.2 to evaluate the sum.

OpenStudy (anonymous):

Theorem 4.2

OpenStudy (anonymous):

Question

OpenStudy (superdavesuper):

seems like u get all the summation formula in ur theorem section, u just have to apply them to ur Q.... OK u need two more eqns: Summation (c*i) = c*Summation (i) where c is a constant; and Summation (f(i) + g(i)) = Summation f(i) + Summation g(i)

OpenStudy (anonymous):

But I don't know which I have to apply to that Q. I don't really understand it

OpenStudy (superdavesuper):

First use this to break up the expression in ur Q: (f(i) + g(i)) = Summation f(i) + Summation g(i)

OpenStudy (anonymous):

so: (2i +31) = Sim. 2i + Sum. 3(i) ?

OpenStudy (superdavesuper):

The expression inside summation is (2i-3) right?

OpenStudy (anonymous):

yes

OpenStudy (superdavesuper):

so using the two eqns i wrote above....u can re-write the prob....

OpenStudy (anonymous):

So what I wrote was wrong?

OpenStudy (superdavesuper):

should be Sum. 2i + Sum.3; not i following the constant 3 right?

OpenStudy (superdavesuper):

Oooops should be 2i-3....so plz re-write the eqn....

OpenStudy (anonymous):

ah ok so: (2i - 3) = Sum. 2i - Sum 3

OpenStudy (superdavesuper):

Correct! now use Summation (c*i) = c*Summation (i) where c is a constant on the first part....

OpenStudy (anonymous):

sum. 2*i = 2 * sum (i)

OpenStudy (superdavesuper):

sum(i) is in ur theorem i believe....

OpenStudy (anonymous):

so does sum(i) become 3(3+1) / 2?

OpenStudy (superdavesuper):

where does the 3 come from??

OpenStudy (anonymous):

hey sorry I had to go. wouldn't (n) be the constant 3? or no?

OpenStudy (superdavesuper):

n is constant but it is a different constant....not 3. plz read ur theorem n the prob itself again n u will see :)

OpenStudy (anonymous):

I did. but there are only two constants...

OpenStudy (superdavesuper):

c and n in the theorem section right?

OpenStudy (anonymous):

therom #1 is c and theorem #2 shows n

OpenStudy (superdavesuper):

ok and in this case, what is n? hint: it is not 3 :)

OpenStudy (anonymous):

2?

OpenStudy (anonymous):

superdave I have to go to sleep will you be around tomorrow to work on this? please?

OpenStudy (superdavesuper):

Sure :) good nite!

OpenStudy (anonymous):

thanks so much!!

OpenStudy (anonymous):

@ganeshie8 @eliassaab could you guys help me please?

ganeshie8 (ganeshie8):

\(\large \sum \limits_{i=1}^{15} (2i -3)\) \(\large \sum \limits_{i=1}^{15} 2i - \sum \limits_{i=1}^{15} 3\) \(\large 2 \sum \limits_{i=1}^{15} i - 3 \sum \limits_{i=1}^{15} 1\) \(\large 2 \frac{15(15+1)}{2} - 3 \times 15\)

ganeshie8 (ganeshie8):

simplify

OpenStudy (anonymous):

uhm what does large stand for?

ganeshie8 (ganeshie8):

oh there must be an issue with latex. u should not see large and all... let me take a snapshot and attach :)

OpenStudy (anonymous):

thank you

ganeshie8 (ganeshie8):

^^ attached wat i see from my side

OpenStudy (anonymous):

hey sorry ... ok so let me simplify

OpenStudy (anonymous):

2 ( 75 + 15 / 2) - 45

OpenStudy (anonymous):

2 (90 / 2) - 45

OpenStudy (anonymous):

2( 45) - 45

OpenStudy (anonymous):

90 - 45 = 45

OpenStudy (anonymous):

45

OpenStudy (anonymous):

@lovablenene I'm done right?

OpenStudy (anonymous):

2* 15(15+1)/2 - 3*15 15(16) - 45 195

OpenStudy (anonymous):

Ok I see now. Thank you